# Assessing Inf luential Observations in Pain Prediction using fMRI Data

### v2

@@ 6 Analysis of Neurologic Signature of Physical Pain
Two metrics are considered to assess detection outcome: $\widehat{\text{I}}^{\text{infl}}_{\text{Detection}(\text{Selector})}$ and $|\widehat{\text{I}}^{\text{infl}}_{\text{Detection}(\text{Selector})}\rvert$. They refer to the estimated index set of influential points and its cardinality, respectively. To obtain them, the existing methods, MIP and DF(LASSO), and the ClusMIP procedure applied to LASSO, SLASSO, ENET, SCAD and MCP are performed. Values for these two metrics for the different detection procedures are reported in Table 1. Since 10-fold CV is used for selecting tuning parameters and every choice is random, the analysis is repeated 1000 times. Since LASSO is used for both DF(LASSO) and ClusMIP(LASSO), to avoid ambiguity, results for both procedures are shown separately in Figures 4 and 5.
@@ 6.1 Detection of Influential Observations
Furthermore, we observe that $\widehat{\text{I}}^{\text{infl}}_{\text{Common}}\coloneqq\ \widehat{\text{I}}^{\text{infl}}_{\text{ClusMIP}(\text{LASSO})}\ \cap\ \widehat{\text{I}}^{\text{infl}}_{\text{ClusMIP}(\text{ENET})}\ \cap\ \widehat{\text{I}}^{\text{infl}}_{\text{ClusMIP}(\text{SCAD})}\ \cap\ \widehat{\text{I}}^{\text{infl}}_{\text{ClusMIP}(\text{MCP})}=\{49,52,55\}$. Hence, observations 49, 52 and 55 are flagged for further investigation. The first two observations correspond to two different temperatures measured on the same participant, indicating this may be a participant that needs to be further investigated. In addition, the first and third observations correspond to the lowest temperature, possibly indicating participants that have a differing pain threshold than other participants. Indeed, a plot of the pain ratings as function of temperature with influential points identified by the ClusMIP(MCP) (Figure 3(A)) indicates that the majority of outliers correspond to low pain ratings at each temperature, a pattern which holds for ClusMIP across all model selectors. For illustration, only the plot for ClusMIP(MCP) is shown (Figure 3(A)) while the other plots in the ClusMIP family are presented in Section 12 of the Supplementary Materials Part I. In fact, the low pain score primarily contributes to the “outlyingness” of the identified influential points, as no anomalies in the feature space are detected as seen in the multi-dimensional scaling plot (Figure 3(C)). This may be explained by the fact that this temperature is generally considered warm, but not always painful, thus differentiating it from other temperature levels. Together these two panels imply that the low pain ratings may need to be described using a different model than that of high pains in fMRI-based thermal pain modeling. This assertion is further strengthened by research on pain processing mechanisms suggesting a discrepancy in brain activity between low and intense pain conditions [9]. In contrast, the outliers of DF(LASSO) do not exhibit an easily recognizable pattern (Figure 3(B)), consisting of both low and high pain scores. This may be due to its limited ability to detect multiple outliers, causing challenges for meaningful interpretation and reliable inference.
@@ 6.2 Effects of Influential Observations on Variable Selection
Figure 4: ClusMIP. (A): average sizes of submodels selected from the full and reduced datasets upon removing $\widehat{\text{I}}^{\text{infl}}_{\text{ClusMIP(Selector)}}$ and $\widehat{\text{I}}^{\text{infl}}_{\text{Common}}$. (B): average number of predictors included in submodels selected from both full and reduced datasets, unique to those from the full dataset, and unique to those from the reduced dataset. (C): change in the normed coefficient estimates upon outlier removal. (D): sample correlation between observed and predicted pain scores. (E): $-\log{(\text{MSE of Prediction})}$.
Figure 5: DF(LASSO). (A): average size of submodels chosen from the full and reduced datasets (removing $\widehat{\text{I}}^{\text{infl}}_{\text{DF(LASSO)}}$). (B): average number of predictors that are in the submodels selected from both full and reduced datasets, unique to those from the full dataset, and unique to those from the reduced dataset. (C): sample correlation between observed and predicted pain scores. (D): $-\log{(\text{MSE of Prediction})}$.
@@ 6.3 Effects of Influential Observations on Coefficient Estimation
In Figure 4(C), we examine the impact of influential points on coefficient estimation by computing the percentage change in the $\ell_{2}$-normed coefficient estimates defined by $(||\widehat{\boldsymbol{\beta}}_{\text{full}}||_{2}-||\widehat{\boldsymbol{\beta}}_{\text{redu}}||_{2})/||\widehat{\boldsymbol{\beta}}_{\text{full}}||_{2}$, where $\widehat{\boldsymbol{\beta}}_{\text{full}}$ and $\widehat{\boldsymbol{\beta}}_{\text{redu}}$ are the coefficient estimates obtained by fitting the full dataset and reduced dataset via removing $\widehat{\text{I}}^{\text{infl}}_{\text{ClusMIP(Selector)}}$ or $\widehat{\text{I}}^{\text{infl}}_{\text{Common}}$. From Figure 4(C), at least $30\%$ change in the $\ell_{2}$-normed coefficients is seen for all model selectors from these two reduced datasets. The largest changes, $63.03\%$ and $82.94\%$, are observed for SCAD. In comparison, although LASSO shows the smallest coefficient adjustments, they reach significant levels of $33.11\%$ and $31.53\%$. The same benchmark is also calculated for DF(LASSO) upon removing $\widehat{\text{I}}^{\text{infl}}_{\text{DF(LASSO)}}$ only, which attains $34.22\%$. These results highlight the considerable impact of outliers on coefficient estimation.
@@ 7 Conclusion and Further Discussion
Using these guiding principles, the GDF measure was proposed in (3), and its exchangeability and CLT established in Theorems 1 and 2. Combination of the GDF and consistent high-dimensional clustering technique leads to our ClusMIP procedure (Algorithm ClusMIP), which controls the FPR in both finite- and large-sample (Theorem 3). Based on the simulations in Section 5, the ClusMIP procedure demonstrates a clear advantage over the existing methods (DF(LASSO) and MIP) in terms of improved detection power, leading to considerably enhanced prediction power (lower MSE in prediction) and variable selection performance (higher empirical probability in selecting a submodel containing the true model) after removal of detected outliers. Specifically, the detection power of ClusMIP(SLASSO) is consistently performs best, varying between 0.90 and 1.00. In contrast, under-performance is seen for DF(LASSO) and MIP, where the detection power reaches levels between 0 and 0.20.

### v3

@@ 6 Analysis of Neurologic Signature of Physical Pain
Two metrics are considered to assess detection outcome: $\widehat{\text{I}}^{\text{infl}}_{\text{Detection}(\text{Selector})}$ and $|\widehat{\text{I}}^{\text{infl}}_{\text{Detection}(\text{Selector})}\rvert$. They refer to the estimated index set of influential points and its cardinality, respectively. To obtain them, the existing methods, MIP and DF(LASSO), and the ClusMIP procedure applied to LASSO, SLASSO, ENET, SCAD and MCP are performed. When computing these two assessment measures, the tuning parameter is chosen as follows: for LASSO, ten-fold CV is used; for SLASSO, the tuning parameter is fixed as $\sqrt{(2\log(p)/n)}$ [29]; for ENET, two-dimensional ten-fold CV is used to select the tuning parameter pair; for SCAD and MCP, the thresholding parameters are set to 3.7 and 3.0, respectively [ncvreg], while the other parameter controlling the scale of shrinkage is determined via ten-fold CV. Here, we recognize that the selection of tuning parameter(s) via $k$-fold CV is stochastic and this randomness may impact the calculation of detected influential points. This necessitates mitigation of the embedded randomness, leading to the detection of more reliable and reproducible outliers. One remedial procedure is repeated $k$-fold CV, where $k$-fold CV is repeated a specified number of times. According to our simulation study provided in Section 13 of the Supplementary Materials Part I (Zhang, Asgharian and Lindquist (2026)), the ClusMIP procedure used with the model selectors under consideration exhibits robustness to data-dependent influential point detection. To be more specific, detection outcomes (measured via the power and FPR of detection, proportion of each individual datum labeled as influential and average number of outliers impacting the selection of true predictors) based on the classic (single) ten-fold CV are almost identical to those obtained when using repeated ten-fold CV. As such, the outliers for all model selectors are obtained using a single ten-fold CV (when applicable) and treated as deterministic for subsequent analysis. Values for these two metrics when using different detection procedures are reported in Table 1.
To quantify the benefits of detecting these influential points, each regularized model selectors are fit to both the full and reduced datasets, where the latter dataset is obtain after removing outliers. Considering the random nature of ten-fold CV in tuning parameter selection (except for SLASSO), in contrast to influential point detection, the downstream analyses on variable selection, coefficient estimation, and prediction based on using both the full and reduced datasets are repeated 1000 times. The empirical averages of these associated benchmarks (defined in Sections 6.2 to 6.4) are reported for comparison purposes. Since LASSO is used for both DF(LASSO) and ClusMIP(LASSO), to avoid ambiguity, results for both procedures are shown separately in Figures 4 and 5.
@@ 6.1 Detection of Influential Observations
Furthermore, we observe that $\widehat{\text{I}}^{\text{infl}}_{\text{Common}}\coloneqq\ \widehat{\text{I}}^{\text{infl}}_{\text{ClusMIP}(\text{LASSO})}\ \cap\ \widehat{\text{I}}^{\text{infl}}_{\text{ClusMIP}(\text{ENET})}\ \cap\ \widehat{\text{I}}^{\text{infl}}_{\text{ClusMIP}(\text{SCAD})}\ \cap\ \widehat{\text{I}}^{\text{infl}}_{\text{ClusMIP}(\text{MCP})}=\{49,52,55\}$. Hence, observations 49, 52 and 55 are flagged for further investigation. The first two observations correspond to two different temperatures measured on the same participant, indicating this may be a participant that needs to be further investigated. In addition, the first and third observations correspond to the lowest temperature, possibly indicating participants that have a differing pain threshold than other participants. Indeed, a plot of the pain ratings as function of temperature with influential points identified by the ClusMIP(MCP) (Figure 3(A)) indicates that the majority of outliers correspond to low pain ratings at each temperature, a pattern which holds for ClusMIP across all model selectors. For illustration, only the plot for ClusMIP(MCP) is shown (Figure 3(A)) while the other plots in the ClusMIP family are presented in Section 12 of the Supplementary Materials Part I (Zhang, Asgharian and Lindquist (2026)). In fact, the low pain score primarily contributes to the “outlyingness” of the identified influential points, as no anomalies in the feature space are detected as seen in the multi-dimensional scaling plot (Figure 3(C)). This may be explained by the fact that this temperature is generally considered warm, but not always painful, thus differentiating it from other temperature levels. Together these two panels imply that the low pain ratings may need to be described using a different model than that of high pains in fMRI-based thermal pain modeling. This assertion is further strengthened by research on pain processing mechanisms suggesting a discrepancy in brain activity between low and intense pain conditions [9]. In contrast, the outliers of DF(LASSO) do not exhibit an easily recognizable pattern (Figure 3(B)), consisting of both low and high pain scores. This may be due to its limited ability to detect multiple outliers, causing challenges for meaningful interpretation and reliable inference.
@@ 6.2 Effects of Influential Observations on Variable Selection
Figure 4: ClusMIP. (A): average sizes of submodels selected from the full and reduced datasets upon removing $\widehat{\text{I}}^{\text{infl}}_{\text{ClusMIP(Selector)}}$ and $\widehat{\text{I}}^{\text{infl}}_{\text{Common}}$. (B): average number of predictors included in submodels selected from both full and reduced datasets, unique to those from the full dataset, and unique to those from the reduced dataset. (C): change in the normed coefficient estimates upon outlier removal. (D): sample correlation between observed and predicted pain scores. (E): $\log{(\text{MSE of Prediction})}$.
Figure 5: DF(LASSO). (A): average size of submodels chosen from the full and reduced datasets (removing $\widehat{\text{I}}^{\text{infl}}_{\text{DF(LASSO)}}$). (B): average number of predictors that are in the submodels selected from both full and reduced datasets, unique to those from the full dataset, and unique to those from the reduced dataset. (C): sample correlation between observed and predicted pain scores. (D): $\log{(\text{MSE of Prediction})}$.
@@ 6.3 Effects of Influential Observations on Coefficient Estimation
[…]
@@ 7 Conclusion and Further Discussion
Using these guiding principles, the GDF measure was proposed in (3), and its exchangeability and CLT established in Theorems 1 and 2. Combination of the GDF and consistent high-dimensional clustering technique leads to our ClusMIP procedure (Algorithm ClusMIP), which controls the FPR in both finite- and large-sample (Theorem 3). Based on the simulations in Section 5, the ClusMIP procedure demonstrates a clear advantage over the existing methods (DF(LASSO) and MIP) in terms of improved detection power, leading to considerably enhanced prediction power (lower MSE in prediction) and variable selection performance (higher empirical probability in selecting a submodel containing the true model) after removal of detected outliers. Specifically, the detection power of ClusMIP(SLASSO) is consistently performs best, varying between 0.90 and 1.00. In contrast, under-performance is seen for DF(LASSO) and MIP, where the detection power reaches levels between 0 and 0.20. Moreover, our simulation studies demonstrate that ClusMIP applied to all model selectors identifies notably more outliers than existing methods, which impacts the selection of true predictors, further highlighting its advantage over the the exist methods.